# Overcoming False Illusions in Real-World Face Restoration with Multi-Modal Guided Diffusion Model

### v1

@@ 4.2 Comparisons with State-of-the-art Methods
MGFR is qualitatively and quantitatively compared with state-of-the-art methods in FR. Notably, the model trained in the initial stage, which is a restoration model solely guided by attribute prompts, already achieves superior visual results. The non-reference prior-based methods selected include PSFRGAN Chen et al. (2021), GPEN Yang et al. (2021a), VQFR Gu et al. (2022), CodeFormer Zhou et al. (2022), DR2 Wang et al. (2023b), and DiffBIR Lin et al. (2023), along with reference prior-based methods ASFFNet Li et al. (2020) and DMDNet Li et al. (2023). Particularly, to ensure contrastive fairness during the inference stage, the description text, containing restricted attributes, is obtained through low-resolution processing. In practical applications, however, users can freely set attribute prompts, enabling more precise and comprehensive guidance. For qualitative results comparing ASFFNet and DMDNet, please refer to the Appendix.
@@
Figure 12: The results of our user study. We randomly select face images under multiple test datasets for user study. Our model achieves excellent recovery quality, which can be further enhanced with high quality reference image and identity information guidance.
Comparison on Synthetic Degradations. Firstly, a quantitative comparison of our model without reference images on the synthetically degraded CelebA-Test dataset is conducted without reference image guidance. According to Table 1, our model achieves the best results on all non-reference metrics, indicative of the superior image quality of the results. Due to space limitation, values of SSIM and PSNR of Table 1 are shown in Section C.1. Additionally, the method’s limitations on full-reference metrics are also noted. This phenomenon, preliminarily demonstrated by experiments in Yu et al. (2024); Jinjin et al. (2020), necessitates a reevaluation of the reference value of indicators like PSNR, SSIM, LPIPS, and the proposal of more effective methods to assess advanced FR methods, particularly as quality improves. More qualitative comparison results of our model can be found in Appendix C. Subsequently, Figure 6 and Figure 1 (a) present a qualitative comparison of the MGFR method applied to the Reface-Test dataset. Even in cases of severe degradation, our method successfully produces highly superior facial details guided by the reference image. In addition, we provide a comparison between our model without reference images and MGFR, with a particular focus on FR tasks involving features like double eyelids, pupil color, and finer facial details, such as wrinkles and moles, which cannot be accurately captured without reference image guidance. This further demonstrates the superiority of utilizing reference image guidance in the FR task. Finally, Table 2 offers quantitative comparison results, indicating that our method significantly surpasses other state-of-the-art methods in perceived quality.
We also conduct a user study with a total of 40 participants, comparing MGFR to other approaches. Participants were asked to select the best quality recovery result from these test techniques for each pair of comparison images, or if no reference image was provided, the result that came closest to the Ground Truth. Figure 12 presents the results, which demonstrate that our method outperforms the state-of-the-art methods in terms of recovery quality. Furthermore, the reconstruction effect can be further enhanced by using the reference image guidance.
@@
Comparison on Real-world Degradations. Additionally, our method was tested on real-world LQ images, which involved collecting degraded face images of publicly available images alongside reference images. The qualitative results, presented in Figure 12, demonstrate that the resulting images possess realistic visual effects with minimal facial illusions. More quantitative and qualitative results are presented in Appendix C.
@@ 4.4 Ablation Study
Additional Information Exchange. Unlike Zhang et al. , we have integrated an additional information flow exchange link (Link-UL) from the U-net model to LCA, and a bidirectional information flow link (Link-LR) between LCA and RCA. Table 4 displays the quantitative test results for the presence of the aforementioned information exchange links. Notably, ‘w/o Link-UL’ refers to results obtained with a single information flow from LCA to U-net model. It is evident that additional information flow exchanges result in improved perceived quality.
@@ 4.5 Limitations and Discussion
Moreover, we believe that allowing excessive free facial attribute control in FR tasks is undesirable due to the potential risk of model abuse. Our model can still recover well without attribute text prompts, thus incorporating additional multimodal information marks our initial effort to enhance recovery performance without increasing task complexity. Regarding the reference image, we disregard its potential to ”edit” the restored results. As with previous studies, we aim for the reference image to further enhance the facial details of the restored face image. In summary, we remain committed to the core objective of the restoration task, ensuring that the output remains faithful to the low-quality input. Moreover, given that individuals typically have multiple reference images, leveraging a broader range of reference information may result in more precise facial detail restoration. Additionally, we acknowledge that high-quality reference images may not always be available, thus the potential of utilizing multiple low-quality reference images for co-guidance is left for future investigation.
@@ 5 Conclusion
We introduce MGFR as a pioneering method in real-world face restoration, at the cutting edge of face image restoration technology, capable of using multi-modal information for guidance to achieve realistic visual effects. Simultaneously, MGFR extends the possibilities of face restoration by controlling text prompts with attributes. The proposed Reface-HQ dataset also offers significant potential for advancing the development of face restoration models based on reference images. As the first multi-modal face image restoration model, MGFR establishes a new benchmark for future technological advancements.
@@ Appendix A Reface-HQ Dataset
To address this challenge, we have created a new real-world dataset named Reface-HQ, as shown in Figure 16. The Reface-HQ dataset encompasses high-definition facial images of celebrities, which have been collected from the Internet. Initially, images with inadequate resolution (minimum 512), low quality and outliers lacking facial features were eliminated. Subsequently, identities represented by fewer than two images were excluded, and face image crop alignment was conducted. Each identity was also manually inspected to eliminate discrepancies in age and makeup. Additionally, to enhance the fairness and inclusiveness of the algorithm, we meticulously review the dataset to ensure it includes samples from all races and skin colors. We strive to ensure the diversity of the training data, thereby minimizing algorithmic bias and discrimination, and further enhancing the algorithm’s fairness and inclusiveness. In summary, Reface-HQ encompasses 5,250 identities, totaling 23500 images with a resolution of 512, subsequently partitioned into three segments: 4870 identities for the training set and 380 for the Reface-Test. The comparison of datasets available for special face restoration tasks is shown in Table 5. IDiff-Face Boutros et al. (2023) is a composite dataset with an indefinite number of images.
@@ C.1 Additional Supplement to Table 1
Appendix D Inference Time and Memory Consuming Analysis
Our method involves two stages of training, and it is clear that our proposed model requires substantial training resources. However, for diffusion-based models, this duration of training is relatively short. We believe the investment is worthwhile as the proposed MGFR model delivers strong results. Additionally, Table 12 displays the average time and memory consumption for model inference. It is important to emphasize that the CFG strategy can be applied to all LDM-based recovery models. Consequently, the results are displayed separately to indicate the use of the CFG strategy during the inference process. Clearly, when our model does not utilize the CFG strategy, time and memory consumption improve only marginally compared to DiffBIR. DR2 boasts the fastest inference time; however, its restoration results are unsatisfactory due to the lack of an additional controller structure. Nevertheless, we assert that efficiency is not the primary focus of this work.
Method Average time (s) Memory consuming (M)
DiffBIR 5.1 11260
DR2 2.6 3144
Ours(w/o CFG) 6.9 15351
Ours (w/ CFG) 12.5 15351
Table 12: Inference Time and Memory Consuming compared with other diffusion model-based methods.

### v2

@@ 4.2 Comparisons with State-of-the-art Methods
MGFR is qualitatively and quantitatively compared with state-of-the-art methods in FR. Notably, the model trained in the initial stage, which is a restoration model solely guided by attribute prompts, already achieves superior visual results. The non-reference prior-based methods selected include PSFRGAN (Chen et al., 2021), GPEN (Yang et al., 2021a), VQFR (Gu et al., 2022), CodeFormer (Zhou et al., 2022), DR2 (Wang et al., 2023b), BFRffusion (Chen et al., 2024) and DiffBIR (Lin et al., 2023), along with reference prior-based methods ASFFNet (Li et al., 2020b) and DMDNet (Li et al., 2023). Particularly, to ensure contrastive fairness during the inference stage, the description text, containing restricted attributes, is obtained through low-resolution processing. In practical applications, however, users can freely set attribute prompts, enabling more precise and comprehensive guidance. For qualitative results comparing ASFFNet and DMDNet, please refer to the Appendix.
@@
Figure 12: The results of our user study. We randomly select face images under multiple test datasets for user study. Our model achieves excellent recovery quality, which can be further enhanced with high-quality reference image and identity information guidance.
Comparison on Synthetic Degradations. Firstly, a quantitative comparison of our model without reference images on the synthetically degraded CelebA-Test dataset is conducted without reference image guidance. According to Table 1, our model achieves the best results on all non-reference metrics, indicative of the superior image quality of the results. Due to space limitation, the values of SSIM and PSNR of Table 1 are shown in Section C.1. Additionally, the method’s limitations on full-reference metrics are also noted. This phenomenon, preliminarily demonstrated by experiments in (Yu et al., 2024; Jinjin et al., 2020), necessitates a reevaluation of the reference value of indicators like PSNR, SSIM, LPIPS, and the proposal of more effective methods to assess advanced FR methods, particularly as quality improves. More qualitative comparison results of our model can be found in Appendix C. Subsequently, Figure 6 and Figure 1 (a) present a qualitative comparison of the MGFR method applied to the Reface-Test dataset. Even in cases of severe degradation, our method successfully produces highly superior facial details guided by the reference image. In addition, we provide a comparison between our model without reference images and MGFR, with a particular focus on FR tasks involving features like double eyelids, pupil color, and finer facial details, such as wrinkles and moles, which cannot be accurately captured without reference image guidance. This further demonstrates the superiority of utilizing reference image guidance in the FR task. Finally, Table 2 offers quantitative comparison results, indicating that our method significantly surpasses other state-of-the-art methods in perceived quality.
We also conducted a user study with a total of 40 participants, comparing MGFR to other approaches. Participants were asked to select the best quality recovery result from these test techniques for each pair of comparison images, or if no reference image was provided, the result that came closest to the Ground Truth. Figure 12 presents the results, which demonstrate that our method outperforms the state-of-the-art methods in terms of recovery quality. Furthermore, the reconstruction effect can be further enhanced by using the reference image guidance.
@@
Comparison on Real-world Degradations. Additionally, our method was tested on real-world LQ images, which involved collecting degraded face images of publicly available images alongside reference images. The qualitative results, presented in Figure 12, demonstrate that the resulting images possess realistic visual effects with minimal facial illusions. More results are presented in Appendix C.
@@ 4.4 Ablation Study
Additional Information Exchange. Unlike (Zhang et al., 2023), we have integrated an additional information flow exchange link (Link-UL) from the U-net model to LCA, and a bidirectional information flow link (Link-LR) between LCA and RCA. Table 4 displays the quantitative test results for the presence of the aforementioned information exchange links. Notably, ‘w/o Link-UL’ refers to results obtained with a single information flow from LCA to U-net model. It is evident that additional information flow exchanges result in improved perceived quality.
@@ 4.5 Limitations and Discussion
Moreover, we believe that excessive flexibility in controlling facial attributes poses a risk of model misuse. Our model demonstrates robust recovery even in the absence of attribute hints. Therefore, for the first time, we incorporate additional multimodal information to enhance recovery quality without increasing task complexity. The reference image serves solely to enhance facial details and does not alter the recovered results. In summary, we remain committed to the core objective of the recovery task: ensuring that the output remains faithful to the low-quality input. Furthermore, leveraging multiple reference images enables more precise detail recovery, while the exploration of utilizing multiple low-quality reference images for collaborative guidance is left for future research.
@@ 5 Conclusion
We introduce MGFR as a pioneering method in real-world face restoration at the cutting edge of face image restoration technology, capable of using multi-modal information for guidance to achieve realistic visual effects. Simultaneously, MGFR extends the possibilities of face restoration by controlling text prompts with attributes. The proposed Reface-HQ dataset also offers significant potential for advancing the development of face restoration models based on reference images. As the first multi-modal face image restoration model, MGFR establishes a new benchmark for future technological advancements.
Acknowledge
This work was supported by the National Key Research and Development Program of China (2024YFB2907500).
@@ Appendix A Reface-HQ Dataset
To address this challenge, we have created a new real-world dataset named Reface-HQ, as shown in Figure 16. The Reface-HQ dataset encompasses high-definition facial images of celebrities, which have been collected from the Internet. Initially, images with inadequate resolution (minimum 512), low quality and outliers lacking facial features were eliminated. Subsequently, identities represented by fewer than two images were excluded, and face image crop alignment was conducted. Each identity was also manually inspected to eliminate discrepancies in age and makeup. Additionally, to enhance the fairness and inclusiveness of the algorithm, we meticulously review the dataset to ensure it includes samples from all races and skin colors. We strive to ensure the diversity of the training data, thereby minimizing algorithmic bias and discrimination, and further enhancing the algorithm’s fairness and inclusiveness. In summary, Reface-HQ encompasses 4,800 identities, totaling 21500 images with a resolution of 512, subsequently partitioned into three segments: 4520 identities for the training set and 280 for the Reface-Test. The comparison of datasets available for special face restoration tasks is shown in Table 5. IDiff-Face Boutros et al. (2023) is a composite dataset with an indefinite number of images.
@@ C.1 Additional Supplement to Table 1
Appendix D Training and Inference Consuming Analysis
In terms of training consumption, the proposed MGFR model employs a two-stage training strategy for the dual-control adapter, leading to a moderate increase in training cost. However, for the diffusion-based image restoration model, this additional training time remains relatively short. Nonetheless, this investment is justified, as the proposed MGFR model demonstrates excellent recovery performance. Additionally, the dual-control adapter’s specialized design enables superior restoration results depend on the guidance of multimodal information. Our experiments (Figure 32) confirm that employing a single traditional adapter structure for multimodal input often results in redundancy between the reference image and the low-quality input, as well as color inconsistencies in the recovered output. This observation, however, does not preclude further exploration in this area. Our future work will focus on employing a specially designed single-transformer adapter to replace the dual-control adapter, aiming to reduce the model’s complexity.
In addition, Table 12 presents the average inference time, memory consumption, parameter count, and FLOPs statistics. Notably, the CFG strategy is compatible with all LDM-based recovery models. Results are presented separately to reflect the CFG strategy’s influence during inference. Without the CFG strategy, our model exhibits slightly higher time and memory consumption compared to DiffBIR Lin et al. (2023). DR2 Wang et al. (2023b) and BFRfusion Chen et al. (2024) exhibit faster inference times; however, their recovery performance is suboptimal. Furthermore, SUPIR’s large-scale model design results in significantly higher training and testing costs compared to other methods, including MGFR. However, MGFR outperforms SUPIR Yu et al. (2024) on the face image restoration task while incurring lower costs (see Appendix M). It should be noted that efficiency is not the primary focus of this work. Moreover, we believe that the development of efficient lightweight models is grounded in the superior performance of large-scale models. Our future iterations will explore model compression techniques, such as quantization and pruning, to enhance inference speed and reduce parameter counts while maintaining MGFR’s superior performance on the face recovery task.
Method Average time (s) Memory consuming (M) #Params (M) FLOPs (G)
DiffBIR 5.1 11260 1716.7 897.5
DR2 2.6 3144 93.56 388.94
BFRffusion 3.2 8338 1197.4 784.5
SUPIR 47.6 54318 3870.0 11950
Ours(w/o CFG) 6.9 15351 2029.3 890.5
Ours (w/ CFG) 12.5 15351 2029.3 2672.4
Table 12: Inference consuming compared with other diffusion model-based methods.
@@ Appendix I More Qualitative Comparisons For MGFR Model
Appendix J Scalability of MGFR for real-world video face restoration
The proposed MGFR framework shows significant potential for real-world video-based face recovery tasks. Unlike single-image restoration, video restoration poses the unique challenge of ensuring temporal consistency. To address this, our method leverages the recovered output of the previous frame as a reference for the current frame. This approach aligns seamlessly with our model architecture, which integrates high-quality continuous frame references into guided restoration. Additionally, as our model does not require strict alignment between the reference and low-quality inputs, it effectively handles natural variations in pose and expression commonly found in consecutive video frames, surpassing previous reference-based face restoration models. By leveraging temporal dependencies between frames, the proposed method ensures identity consistency and high-quality recovery in video sequences. Future work could enhance this approach by integrating explicit temporal models or constraints, such as optical flow guidance, to better handle motion artifacts and dynamic variations in video data. Unlike single-image restoration based on reference images, video data offers more diverse and abundant training samples, which we believe will further unlock the potential of our proposed model. This will be a key focus of our future work.
@@
Appendix K Model stability
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption LR Seed 2982 Seed 78 seed 11281 seed 15 seed 424 seed 9168 GT
Figure 29: Model Stability Analysis. The recovery results of MGFR remain consistent across different random seeds, eliminating the need for selection among multiple input outcomes.
Appendix L Brief Overview of Evaluation Metrics
For quantitative comparison, the selected image quality evaluation metrics include full-reference metrics PSNR, SSIM, and LPIPS Zhang et al. (2018). Yu et al. (2024); Jinjin et al. (2020) experiment initially confirmed that as image restoration quality improves, the reference utility of metrics such as PSNR, SSIM, and LPIPS needs to be re-evaluated, necessitating the selection of more effective evaluation indicators. Therefore, we introduce three non-reference metrics—ManIQA Yang et al. (2022), ClipIQA Wang et al. (2023a), and MUSIQ Ke et al. (2021)—in this work.
A summary of each evaluation metric is provided below.
  • SSIM is a key metric for assessing image restoration quality, measuring the similarity between the restored and original images based on brightness, contrast, and structural information. It has been widely used in previous face image restoration tasks Lin et al. (2023); Wang et al. (2023b; 2021a); Yang et al. (2021a); Zhou et al. (2022); Gu et al. (2022); Yu et al. (2024); Chan et al. (2021); Chen et al. (2021); Li et al. (2023; 2020b; 2020a); Wang et al. (2021b); Li et al. (2020c); Teng et al. (2022).
  • PSNR is a metric derived from the mean square error (MSE), calculated as the logarithmic ratio of the maximum possible pixel value to the error. The results are expressed in decibels (dB), where higher values signify better image quality. It has been widely used in previous face image restoration tasks Lin et al. (2023); Wang et al. (2023b; 2021a); Yang et al. (2021a); Zhou et al. (2022); Gu et al. (2022); Dogan et al. (2019); Yu et al. (2024); Chan et al. (2021); Chen et al. (2021); Li et al. (2023; 2020b; 2020a); Wang et al. (2021b); Li et al. (2020c); Teng et al. (2022).
  • LPIPS quantifies image differences by extracting features from deep neural networks and measuring the distances between these features. This metric better captures perceptual changes in image details and textures. Previous studies have emphasized image similarity metrics aligned with human visual perception Lin et al. (2023); Wang et al. (2023b; 2021a); Yang et al. (2021a); Zhou et al. (2022); Gu et al. (2022); Yu et al. (2024); Chan et al. (2021); Chen et al. (2021); Li et al. (2023; 2020b; 2020a); Wang et al. (2021b); Li et al. (2020c); Teng et al. (2022).
  • ManIQA maps images into a low-dimensional manifold space and analyzes their feature distribution and location to assess image quality. This approach demonstrates a high correlation with perceived quality, and its effectiveness has been validated in Yu et al. (2024).
  • MUSIQ implements a multi-scale feature extraction mechanism designed to capture the quality characteristics of images across varying resolutions and perceptual scales for effective image quality evaluation, and its effectiveness has been validated in Yu et al. (2024).
  • ClipIQA leverages the robust vision-language priors embedded within the CLIP model. The focus is on enhancing the capability to evaluate both quality perception (seeing) and abstract perception (feeling) of visual content. This approach’s effectiveness has been demonstrated in Yu et al. (2024).
Appendix M Qualitative Comparison with SUPIR
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption LR SUPIR Ours w/o Ref. GT
Figure 30: Qualitative comparisons with SUPIR Yu et al. (2024) for our text-guided baseline model on synthetic dataset under moderate degradation in CelebA-Test dataset. Zoom in for the best view.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption LR SUPIR Ours w/ Reference GT Reference
Figure 31: Qualitative comparisons with SUPIR Yu et al. (2024) for MGFR on synthetic dataset under moderate degradation in Reface-Test dataset. Zoom in for best view.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption LR Ours w/ SCA Ours w/ DCA GT Reference
Figure 32: Ablation experiments comparing the reception of multi-modal information using a single control adapter (SCA) versus a dual control adapter (DCA) revealed that SCA led to reduced recovery performance and increased chromatic aberration.